### v1

# Waste Heat and Habitability: Constraints from Technological Energy Consumption

Abstract
Waste heat production represents an inevitable consequence of energy conversion as per the laws of thermodynamics. Based on this fact, by using simple theoretical models, we analyze constraints on the habitability of Earth-like terrestrial planets hosting putative technological species and technospheres characterized by persistent exponential growth of energy consumption and waste heat generation: in particular, we quantify the deleterious effects of rising surface temperature on biospheric processes and the eventual loss of liquid water. Irrespective of whether these sources of energy are ultimately stellar or planetary (e.g., nuclear, fossil fuels) in nature, we demonstrate that the loss of habitable conditions on such terrestrial planets may be expected to occur on timescales of $\lesssim 1000$ years, as measured from the start of the exponential phase, provided that the annual growth rate of energy consumption is of order $1\%$. We conclude by discussing the types of evolutionary trajectories that might be feasible for industrialized technological species, and sketch the ensuing implications for technosignature searches.
keywords:
astrobiology – Earth (planet) – habitable planets – habitable zone – planetary climates – technosignatures
1 Introduction
It is now firmly established that the upsurge in atmospheric CO2 concentration caused by human technological activities has been elevating Earth’s average surface temperature (e.g., Archer, 2016; Hartmann, 2016; Dessler, 2021; Magnan et al., 2021; Abbass et al., 2022; Romm, 2022; Tziperman, 2022). Nevertheless, it is worth appreciating that ongoing anthropogenic warming is not exclusively a consequence of the intensified greenhouse effect stemming from fossil fuel utilization (though the latter phenomenon is unequivocally proven to be the chief driver). Instead, as elucidated in a number of analyses and overviews (e.g., Chaisson, 2008; Flanner, 2009; Kleidon, 2012; Ming et al., 2014; Lu et al., 2017; Kleidon et al., 2023), this facet embodies a broader outcome of energy production and conversion underpinning modern (viz., globalized and industrialized) human societies. The pioneering 1969 publication by the Soviet scientist Mikhail I. Budyko explicitly underscored this theme (Budyko, 1969, pg. 618):
All the energy used by man is transformed into heat, the main portion of this energy being an additional source of heat as compared to the present radiation gain. Simple calculations show (Budyko 1961) that with the present rate of growth of using energy the heat produced by man in less than two hundred years will be comparable with the energy coming from the sun.
This feature arises from the fundamental fact that energy transduction inevitably generates waste heat (e.g., Budyko, 1969; Summers, 1971; Smil, 2008) [^1] Interestingly, around the same time as Budyko’s paper, Larry Niven’s classic science fiction novel Ringworld (1970) also acknowledged the importance of waste heat as an evolutionary factor for technological species (Niven, 1970).. Although the current contribution of this factor to global warming on Earth is negligible compared to the impact of carbon emissions (Ming et al., 2014; Kleidon et al., 2023), [^2] In localized regions endowed with high industrialization (e.g., large and dense cities), however, it is already a notable contributor to warming (Ichinose et al., 1999; Block et al., 2004; Ohashi et al., 2007; Allen et al., 2011; Zhang et al., 2013; Bohnenstengel et al., 2014; Salamanca et al., 2014; Stewart and Kennedy, 2017; Sun et al., 2018; Raj et al., 2020; Molnár et al., 2020), and this effect is anticipated to increase in the future. it is expected to amplify quite substantially in the next century in the event that the current growth rate of global energy consumption persists (Lu et al., 2017); the consequent rise in global temperature (just from this process) derived from assorted climate models may reach up to $\sim 1$ K in this duration (Flanner, 2009; Nickolaenko, 2009; Steiglechner et al., 2019).
Cruder estimates, when extrapolated over longer intervals, project a temperature increase of $\sim 10$ K over a period of $350$ years, upon assuming similar energy trends as today (Chaisson, 2008). Likewise, a cognate analysis by Mullan and Haqq-Misra (2019) (see also von Hoerner 1975; Dutil and Dumas 2009) concluded that direct heating engendered by progressively higher energy consumption consistent with current rates could dramatically impede human habitability on timescales of merely hundreds of years in the future. It is crucial to realize that this contribution to global warming would persist even in the scenario of achieving a complete transition to zero CO2-emissions energy production.
While this effect is patently of profound importance for the long-term future and fate of human civilization(s), it also has broader significance in the general context of planetary evolution and the potential for harboring technologically capable life in the universe – this latter domain is closely linked with detecting technosignatures, a burgeoning area of astrobiology (Ćirković, 2012; Lingam and Loeb, 2021; Socas-Navarro et al., 2021), with roots traceable to the 1960s (Sagan, 1963; Drake, 1965; Shklovskii and Sagan, 1966; Tarter, 2001). To be specific, we are free to inquire whether the energetic growth, often innately presumed to be associated with technological activity (Kardashev, 1985), imposes a universal fundamental limit on the habitability of a planetary environment, arising simply from thermodynamic considerations, as indicated in von Hoerner (1975).
The notion that technological advancement is correlated with the amount of energy consumed was initially articulated vis-à-vis extraterrestrial life by Kardashev (1964); early studies implicitly adopted the premise that technological species would pursue and evince growth of their (energetic) activities (Dyson, 1960; Shklovskii and Sagan, 1966; Kardashev, 1985). Kardashev’s idea has been predominantly employed to categorize hypothetical extraterrestrial technological entities, [^3] We remark, however, that a few proposed alternatives to the Kardashev scale exist (Sagan, 1973; Zubrin, 2000; Galantai, 2005; Frank et al., 2017; Ivanov et al., 2020; Lingam et al., 2023). and gauge their detectability through the types of technosignatures they might produce (Ćirković, 2015; Gray, 2020; Lingam and Loeb, 2021), but studies that strive to understand how heightened energy consumption by these entities impacts a planetary environment and whether this trend actually aligns with long-term habitability are comparatively limited in scope and in number.
Thus, the summum bonum of this paper is to perform a preliminary investigation of how the habitability of Earth-like planets would be affected by the presence of life specifically capable of engaging in progressively energy-demanding technological activities. One of our primary objectives in this vein is to gauge the maximum lifespan of habitable planets that harbor an evolving “technosphere” (Donges et al., 2017; Zalasiewicz et al., 2017; Herrmann-Pillath, 2018), by making certain assumptions regarding the growth of energy consumption by an industrialized technological species. This analysis parallels the canonical approach used to estimate the lifetime of continuously habitable zones via incorporation of the increase in stellar luminosity, which exclusively accounts for abiotic factors (see, e.g., Kasting et al., 1993; Rushby et al., 2013; Gallet et al., 2017; Tuchow and Wright, 2023), and not biological/technological feedbacks.
The outline of the paper is as follows. We present the methods and accompanying results pertaining to the impacts of global heating (from energy production) on putative biospheres (and technospheres) and climates in Section 2, where it is shown that the loss of habitable conditions might occur in $\sim 10^{3}$ yr. Next, we elucidate their ramifications for the potential trajectories of technological species (e.g., collapse; slowdown or cessation of growth; expansion into space) and detecting technosignatures in Section 3.
2 Methods and Results
In this Section, we describe some of the thermodynamic principles underlying our modeling, and delineate the results that ensue in our framework.
2.1 Thermodynamics of energy production
The Second Law of thermodynamics dictates that whenever energy $E$ is converted into useful work $W$, waste heat $\mathcal{Q}$ is produced and dissipated in the environment (e.g., Summers, 1971; Struchtrup, 2024; Blundell and Blundell, 2010; Schroeder, 2020). If the conversion mechanism exhibits an efficiency of $\eta=W/E$, the amount of energy that goes into heating is expressible as
$\mathcal{Q}=\left(1-\eta\right)E.$ (1)
Carnot’s theorem canonically sets a theoretical limit on the maximum efficiency attainable by a cyclical process operating between temperatures $T_{C}$ and $T_{H}$ (e.g., Struchtrup, 2024; Schroeder, 2020), as given by the famous formula:
$\eta=1-\frac{T_{C}}{T_{H}},$ (2)
with the accompanying condition that $T_{C}\leq T_{H}$.
It is, therefore, unavoidable that energy conversion (to thermodynamic work) enabling technological activity produces a finite amount of waste heat, $\mathcal{Q}_{\rm tech}$ (Budyko, 1969, 1972). When energy is extracted from sources that would otherwise be rendered “inactive” (such as burning fossil fuels or exploiting nuclear reactions), $\mathcal{Q}_{\rm tech}$ will constitute an extra heat source that is fed into the planetary system, in addition to the “natural” heat deposited by stellar irradiation or by the dissipation of “intrinsic” (e.g., radioactivity) planetary energy sources. It is worth recognizing that the warming connected to $\mathcal{Q}_{\rm tech}$ is conceptually different from that caused by the enhanced (viz., anthropogenic) greenhouse effect resulting from alterations of atmospheric chemistry. While in principle a “zero emissions” energy conversion strategy ought to be viable, the production of waste heat can only be minimized, but never completely eliminated.
Although the amount of waste heat dissipated by technological activity is evidently not $100\%$ of the energy input, in most practical applications of current technology, it is comparable to this upper bound. Moreover, waste heat does not seem reducible to a negligible fraction, although it might be curtailed drastically in theory (cf. Bell, 2008; Cullen and Allwood, 2010; Cullen et al., 2011; Jouhara et al., 2018; Garofalo et al., 2020; Geffroy et al., 2021). The energy conversion efficiency attributed to fossil fuels, nuclear and even geothermal sources is distinctly lower than the theoretical Carnot limit. In global society, it has consequently been estimated that $\mathcal{Q}/E$ is typically $\gtrsim 50\%$ (Graus et al., 2007; Smil, 2008, 2017; Cullen and Allwood, 2010; Zarrouk and Moon, 2014; Forman et al., 2016; Firth et al., 2019; Garofalo et al., 2020; Obara and Tanaka, 2021). Renewable energy sources such as wind, hydroelectric, and tidal power may not entail direct additional heating (although they could impact the climate indirectly), but whether their exclusive usage could meet the demands of a global (industrialized) technological society embarking on sustained exponential growth of the type studied herein remains an open question (cf. MacKay, 2009; Smil, 2015, 2020). Finally, the conversion of abundant stellar energy warrants a separate treatment that is tackled in Section 2.3.
As underscored in the preceding exposition, the waste heat is comparable to the input energy for known sources. Therefore, we will hereafter adopt the ostensibly reasonable premise that the waste heat generated is roughly equal to the energy consumption associated with technological activity, unless indicated otherwise. If we postulate that the magnitude of waste heat generated is diminished by about an order of magnitude, this does not alter our qualitative findings, and it modifies the quantitative results by merely a modest degree; this facet is implicit in the paragraph below (17).
2.2 Global heating from technological activity
At equilibrium, the incoming stellar energy reaching a planetary surface has to equal the outgoing flux emitted by the planet itself. The basic (zero-dimensional and equilibrium) energy balance equation can be expressed as (e.g., Jacob 1999, Chapter 7, Vallis 2012, Chapter 1, and Tziperman 2022, Chapter 2):
$\frac{S}{4}(1-A)+\frac{1}{2}\varepsilon\sigma T^{4}+Q_{\rm tech}=\sigma T^{4},$ (3)
where $S$ is the stellar flux incident at the top of the atmosphere, $A$ is the planetary albedo, $\varepsilon$ is the fraction of terrestrial infrared radiation absorbed by the atmosphere (a parameter that can be calibrated to model greenhouse warming), $T$ is the planetary surface temperature (which is not the same as the planet’s equilibrium blackbody temperature), and $Q_{\rm tech}$ is the waste heat flux (units of power per unit area) ensuing from technological activity. Where needed (i.e., not for every calculation), we will henceforth adopt an average albedo of $A\approx 0.3$ for modern Earth (Trenberth et al., 2009; Stephens et al., 2015), as well as for “Earth-like” habitable worlds (i.e., Earth-analogs) (cf. Del Genio et al., 2019; Rushby et al., 2019; Madden and Kaltenegger, 2020). [^4] However, as long as $A$ is not very close to unity (i.e., $A\rightarrow 1$), our forthcoming quantitative results are not particularly sensitive to this parameter.
It should, however, be noted that (3) neglects the supplemental role of convection (Hartmann, 2016; Tziperman, 2022), which may be of importance for certain temperate planets orbiting M-dwarfs (Shields et al., 2016). Likewise, we neglect internal heat sources in (3), which would enter as additional terms on the LHS, and would serve to elevate the temperature further, and might possibly diminish the timescales computed later in this paper. The possible sources are radioactivity and tidal heating, both of which are relatively modest on Earth compared to $S$; the former translates to a heat flux of $\sim 7.4\times 10^{-2}$ Wm-2 (The Borexino Collaboration, 2020). On some habitable worlds, however, it is plausible that tidal heating (Jackson et al., 2008; Barnes et al., 2009; Driscoll and Barnes, 2015; Dobos et al., 2019) and radiogenic heating (Frank et al., 2014; Lingam and Loeb, 2020; Nimmo et al., 2020) could be higher, and thus necessitate inclusion in the global energy budget equation.
The energy balance in (3) is applied at the planetary surface. The first and the second terms on the LHS account for incoming stellar radiation and the thermal radiation emitted by the atmosphere (toward the surface), whereas the RHS is the outgoing thermal radiation from the planetary surface. The last term on the LHS is the extra contribution arising from technological activity. For the sake of the subsequent exposition, we rewrite (3) in terms of the parameter $\tau\equiv 1-\varepsilon/2$, which has been interpreted as an effective “transmittance” in some publications (e.g., Steiglechner, 2018), thereupon simplifying to yield
$\frac{S}{4}(1-A)+Q_{\rm tech}=\tau\sigma T^{4},$ (4)
where it should be noted that $\tau$ is innately dependent on atmospheric properties, such as its composition, and encapsulates a fair amount of complexity.
All values should be inherently perceived as global averages. As mentioned in Section 2.1, $Q_{\rm tech}$ is anticipated to be comparable to the total power converted to support technological activity. As a reference, adopting a global primary energy consumption of $1.8\times 10^{5}$ TWh in 2022 (Ritchie et al., 2024), [^5] https://ourworldindata.org/energy-production-consumption and averaging over the entire surface of the Earth ($5.1\times 10^{14}\ \rm m^{2}$) yields an estimate of $Q_{\rm tech}\sim 0.04$ Wm-2, which matches the calculation reported in Kleidon et al. (2023), acting as a consistency check. This is currently a negligible fraction of the radiative forcing attributable to the anthropogenic greenhouse effect (computed to be $2.7$ Wm-2) responsible for global warming (Kleidon et al., 2023).
The planetary temperature can be further decomposed as $T=T_{0}+\Delta T$, where we have defined
$T_{0}=\left[\frac{S(1-A)}{4\tau\sigma}\right]^{1/4}$ (5)
as the temperature when $Q_{\rm tech}\approx 0$, and $\Delta T$ is the ensuing warming when $Q_{\rm tech}\neq 0$. On solving (4), we find
$\Delta T=\left(T_{0}^{4}+\frac{Q_{\rm tech}}{\tau\sigma}\right)^{1/4}-T_{0}.$ (6)
This expression can also be recast in terms of a fractional temperature increase by dividing throughout with $T_{0}$ as follows:
$\frac{\Delta T}{T_{0}}=\left(1+\frac{4}{1-A}\frac{Q_{\rm tech}}{S}\right)^{1/4}-1,$ (7)
where we have substituted (5) into (6). An interesting feature of the above equation is that the lower bound on $\Delta T/T_{0}$ is attained in the low-albedo limit of $A\rightarrow 0$, and is given by
$\frac{\Delta T}{T_{0}}\geq\left(1+4\frac{Q_{\rm tech}}{S}\right)^{1/4}-1,$ (8)
which is independent of the atmospheric parameter $\tau$, just like the parent equation (7) from which it was derived.
Next, if we suppose that $Q_{\rm tech}/S\ll 1$ in (7), we obtain a simple relation through a Taylor series expansion,
$\frac{\Delta T}{T_{0}}\sim\frac{1}{1-A}\frac{Q_{\rm tech}}{S},$ (9)
which implies that the increase in global surface temperature is linearly proportional to the extent of waste heat generation (encapsulated in $Q_{\rm tech}$); to put it differently, the surface temperature responds in a roughly linear fashion to an increase in energy consumption. The exact same scaling law was predicted by Budyko (1972), thus acting as a consistency check, and this trend remains approximately valid $\sim 50$ years after the paper was published (Lapenis, 2020).
[…]
It must be recalled that our goal in this work is not to perform state-of-the-art climate modeling, but to formulate a heuristic treatment that captures the salient trends (in consideration of the many variables at play). Furthermore, as intimated in Section 2.1 and particularly Section 2.4, the supposition of exponential growth in power consumption and waste heat generation ensures that most of our predictions pertaining to the timescales (of various variables) are quite robust, owing to a weak (i.e., logarithmic) dependence.
2.3 Global heating from stellar energy conversion
Hitherto, $Q_{\rm tech}$ has been interpreted as extra heating, in addition to that naturally produced by stellar radiation, from a source that would otherwise be sequestered within the planet (e.g. by fossil fuel burning or nuclear reactions). We will now analyze the case wherein energy for technological activity is entirely produced via suitable (e.g., photovoltaic) conversion of the stellar flux reaching the planetary surface (refer to De Vos 2008; Weinstein et al. 2015; Khan and Arsalan 2016; Kabir et al. 2018; Gong et al. 2019; Hayat et al. 2019).
Ultimately, an appropriate fraction of such energy converted into useful work (for instance, through the production of electricity) is dissipated as heat. At first, it may appear that this energy usage ought not translate to an elevation of the planetary surface temperature, since all the incoming and outgoing energy is already seemingly included in (4) with $Q_{\rm tech}=0$. However, the efficient conversion of stellar energy necessitates the collection of incident flux that would otherwise be reflected to space by tapping into methods such as surface texturing (Narasimhan and Cui, 2013; Kim et al., 2020b) and anti-reflective coatings (Raut et al., 2011; Sarkın et al., 2020; Shanmugam et al., 2020), which thereby can induce a net decrease of the planetary albedo (Berg et al., 2015; Barron-Gafford et al., 2016; Broadbent et al., 2019; Ziar et al., 2019; Sailor et al., 2021; Xu et al., 2024). [^7] We caution, however, that solar energy conversion by means of photovoltaic cells engenders an intricate cascade of climatic effects on Earth currently, which may contribute to either regional heating (Barron-Gafford et al., 2016; Broadbent et al., 2019; Sailor et al., 2021; Sun et al., 2022; Khan and Santamouris, 2023) or cooling (Masson et al., 2014; Hu et al., 2016; Zhang and Xu, 2020; Xu et al., 2024); by the same token, future extrapolations to extensive deployment of these cells could have major attendant uncertainties. If $f_{s}$ is the fractional area covered by solar panels and $A_{s}$ is the panel albedo, the balance equation becomes:
$\frac{S}{4}\left(1-(1-f_{s})A-f_{s}A_{s}\right)-\tau\sigma T^{4}=0,$ (10)
which reduces to (4) with $Q_{\rm tech}=0$ in the limit of $f_{s}\rightarrow 0$ (i.e., when solar panels and other energy sources like fossil fuels are absent), along expected lines. The above expression is therefore rewritten to highlight this aspect:
$\frac{S}{4}(1-A)+Q_{s}=\tau\sigma T^{4},$ (11)
where $Q_{s}$ constitutes an additional heating term stemming from the decrease in albedo attributable to the collecting area covered by photovoltaic panels:
$Q_{s}\equiv\frac{S}{4}f_{s}\left(A-A_{s}\right).$ (12)
The above result may be viewed as a lower bound because it only comprises the heating from change(s) in albedo, and does not incorporate the finite efficiency of photovoltaics (Kirchartz and Rau, 2018; Kim et al., 2020a; Green et al., 2023), [^8] For a single junction photovoltaic cell, Shockley and Queisser (1961) derived a theoretical efficiency limit, which has been subsequently modified both empirically and theoretically (Nelson, 2003; Krogstrup et al., 2013; Rühle, 2016; Ehrler et al., 2020; Markvart, 2022). and further losses during energy conversion. The fraction of stellar flux converted into usable energy flux $\mathcal{P}$ (e.g., by way of photovoltaics) is
$\mathcal{P}=\frac{S}{4}f_{s}\left(1-A_{s}\right).$ (13)
The corresponding temperature increase can be computed by substituting $Q_{s}$ in (6). By combining the preceding two equations, we find that
$Q_{s}=\left(\frac{A-A_{s}}{1-A_{s}}\right)\mathcal{P}.$ (14)
Hence, the power derived from stellar radiation conversion is responsible for reduced global heating compared to the previous scenario of $Q_{\rm tech}\sim\mathcal{P}$ motivated in Section 2.1 for the category of intrinsic technological power sources (e.g., fossil fuels and nuclear reactions). This statement is a consequence of the fact that the term preceding $\mathcal{P}$ in (14) is smaller than unity.
In the idealized limit of $A_{s}=0$, the contribution to heating can decrease by a factor of $A$ for the same energy consumption (i.e., a fixed value of $\mathcal{P}$), as revealed from inspecting (14); this case is consistent with the simplified results presented in Berg et al. (2015, Section 4). The maximum power attainable via photovoltaic cells (assuming an unrealistic 100% efficiency in stellar energy conversion) corresponds to the case with $f_{s}=1$ and $A_{s}=0$ in (13), to wit, when all the stellar flux reaching the entire planetary surface ($S/4$) is collected by panels, without any reflection to space or miscellaneous losses. In this situation, the planet is anticipated to experience heating by an extra $A\times(S/4)$, as revealed by (14). On plugging these values in (11), it can be shown that $T$ for this planet corresponds to that of a perfect absorber, as expected.
To exceed the aforementioned power limit exclusively by means of stellar power, additional stellar flux would need to be redirected towards the surface, or the collecting area should be increased; potential avenues include the deployment of mirrors or photovoltaic panels in space and the construction of Stapledon-Dyson spheres (e.g., Stapledon, 1937; Dyson, 1960; Korpela et al., 2015). In such scenarios, the outcome is predicted to yield a further temperature increase, which would be effectively regulated by the analog of (6) that is pertinent when external energy is fed into the system.
2.4 Connection to the Kardashev scale
Figure 1: Growth of power consumption $P(t)$ of technological species as per (16), upon assuming a constant annual growth rate $a$ of 0.01 (blue), 0.02 (red) and 0.03 (yellow). The start time corresponds to that of humankind circa 1800 (i.e., cessation of the preindustrial phase), with an associated power consumption of $P_{0}$. The green curve represent the actual data for humanity’s global power consumption (from (Ritchie et al., 2024)). The right axis roughly shows the ensuing global heating flux $Q_{\rm tech}$, obtained from dividing the global power consumption by the total surface area of Earth.
Circling back to the discussion in Section 1, the classic paper by Kardashev (1964) proposed that energy consumption can be used to broadly categorize the technological level of extraterrestrial intelligences. Quantitatively, this classification is expressed using the so-called Kardashev scale, which is mathematically represented as (Gray, 2020):
$K=\frac{1}{10}\log_{10}\left(\frac{P}{10^{6}\,\mathrm{W}}\right)$ (15)
where $P$ is the power available to the technological species and $K$ is its corresponding level in this scale.
A frequently adopted prescription for estimating the time required for an advancing technological species to ascend the rungs of the Kardashev scale is to employ a constant exponential growth rate $a$ of energy utilization, [^9] It is known, however, that presupposing any particular growth pattern involves attendant uncertainty (e.g., Smil, 2000), and may not be tenable, as partly elaborated in Section 3. reflecting the observed trend on Earth during the industrial era (von Hoerner, 1975). This assumption can be quantitatively encapsulated as follows:
$P(t)\approx P_{0}(1+a)^{t/{\rm 1yr}}$ (16)
where $P_{0}$ represents the energy employed by the technological species at some initial time: a suitable choice might be the epoch of transition from pre-industrial to industrial societies, although demarcating such an epoch naturally entails a bevy of accompanying subtleties. If we take the logarithm of (16) and postulate that $a\ll 1$ is valid, we arrive at
$t\,({\rm in\,yr})\approx\frac{1}{a}\ln\left(\frac{P(t)}{P_{0}}\right),$ (17)
on simplification, which reveals that $t$ has a modest dependence on $a$ (inverse proportionality) and a very weak (i.e., logarithmic) dependence on $P_{0}$. Even if the latter were to vary by many orders of magnitude, it ought to alter the predicted value of $t$ by a relatively moderate amount (e.g., an order of magnitude), thus conferring some degree of robustness. These two vital points should be borne in mind with respect to upcoming results. For instance, as remarked toward the ending of Section 2.1, reducing the waste heat production – which is crudely tantamount to diminishing the magnitude of $P$ – ought not modify our conclusions substantively.
For reference, global primary energy consumption at the onset of industrial revolution on Earth (circa 1800) was $P_{0}\approx 6000$ TWh (Smil, 2017), [^10] https://ourworldindata.org/grapher/global-primary-energy and the average growth rate has been $a\sim 0.02$ during the industrial epoch (Ritchie et al., 2024) [^11] https://www.oecd-ilibrary.org/energy/primary-energy-supply/indicator/english_1b33c15a-en. Figure 1 depicts the evolution of $P$ from (16), with various assumptions for the growth rate and some useful values for reference. A consequence of the postulated exponential growth is that the relevant parameter (energy consumption) grows by orders of magnitude on timescales of merely hundreds of years, which is consistent with prior publications centered on this theme (von Hoerner, 1975; Dutil and Dumas, 2009; Haqq-Misra and Baum, 2009; Mullan and Haqq-Misra, 2019; Kopparapu et al., 2024).
While the adoption of a constant growth rate may create the impression that a swift transition to ascending levels of the Kardashev scale is bound to happen, our preceding discussion warrants caution. Specifically, we must take into account the inevitable global heating associated with energy consumption (represented by $P$). If, as assumed implicitly in this paper (see Section 2.1), the waste heat flux $Q_{\rm tech}$ broadly obeys the relation,
$Q_{\rm tech}\sim\frac{P}{4\pi R_{\oplus}^{2}},$ (18)
for Earth-sized worlds, it is anticipated that the concomitant warming would escalate as rapidly as $P$. If $P$ is instead generated through the conversion of stellar flux, the accompanying heating only decreases by a fairly modest factor of $A$ at most, as explained in Section 2.3. However, the maximum power attainable in this scenario is $P=\left(4\pi R_{\oplus}^{2}\right)\times\left(S/4\right)$ (refer to Section 2.3). On substituting $S=S_{\oplus}\equiv 1360$ Wm-2 in this expression for $P$, and using (15), we end up with $K\approx 1.12$. As a result, even the application of perfectly efficient conversion of incident stellar power cannot enable a putative technological species to exceed $K=1$ by a sizable margin.
In the forthcoming Section 2.5, we will further delve into the implications for planetary habitability, and follow this up by calculating the maximal lifetime and Kardashev scale of technological species in Section 2.6.
2.5 Implications of heating for habitability
Figure 2: The fractional increase in planetary temperature when power consumption $P$ grows at a constant annual rate $a$ of 0.01 (blue), 0.02 (red) and 0.03 (yellow). Timescale shown on the horizontal axis corresponds to that of Figure 1. The shaded regions encompass the upper and lower limits for $\Delta T/T_{0}$, with optimistic estimates corresponding to the production of all power via conversion of stellar flux. In plotting the curves, we have assumed a fiducial planetary albedo of $A=0.3$ for Earth-like worlds.
We can apply the previous considerations to estimate the ramifications for planetary habitability ensuing from the production of waste heat associated with energy consumption by technological species. In essence, we are interested in forecasting the time it would take for a planet hosting a technological species – pursuing an exponential growth of energy consumption – to experience a rise in temperature so significant to jeopardize its habitability. This matter was studied by the likes of von Hoerner (1975) and Mullan and Haqq-Misra (2019).
In Figure 2, the fractional temperature increase $\Delta T/T_{0}$ is plotted as a function of time lapsed for different annual growth rates of energy assumption by invoking (7), where the heating term is specified appropriately based on the power source (viz., non-stellar vs stellar energy). The two horizontal lines of $\Delta T/T_{0}=0.02$ and $\Delta T/T_{0}=0.15$ represent the crossing of conceivably vital planetary thresholds for habitability justified below. [^12] These thresholds are partially reminiscent of the central “planetary boundaries” and “planetary climate tipping points” concepts in climate science (Lenton et al., 2008; Rockström et al., 2009; Barnosky et al., 2012; Steffen et al., 2015; Raworth, 2017; O’Neill et al., 2018; Ritchie et al., 2021; Persson et al., 2022; Richardson et al., 2023; Rockström et al., 2023; Wunderling et al., 2024), as well as cognate paradigms, but are not equivalent to them. We notice that a major elevation of surface temperature is expected on short timescales of $\lesssim 1000$ years.
First, let us consider Earth’s biosphere as a reference point. It has been known for well over a century that heat stress poses severe impediments to human activities (Buzan and Huber, 2020), owing to which formulations like the wet-bulb temperature have been introduced for quantifying such effects (Haldane, 1905). Based on heat stress considerations, Sherwood and Huber (2010) proposed that $\Delta T=12$ K could render large swathes of Earth uninhabitable for Placentalia (a taxon which includes humans), which is also consistent with the study by Asseng et al. (2021), who concluded that $\Delta T>10$ K would induce heat stress across many species. Thus, this numerical value was consequently employed by Mullan and Haqq-Misra (2019) in their analysis.
However, it is plausible that $\Delta T=12$ K may not represent an accurate upper bound because a potentially crucial factor could be missing, namely, the intricate interplay of ecological interactions. By incorporating such kinds of dependencies into a suite of numerical models and modeling possible trajectories, Strona and Bradshaw (2018) concluded that a swift temperature increase of merely $\Delta T\approx 6$ K might suffice to drive global biodiversity collapse (see also Song et al. 2021; Malanoski et al. 2024 for elaborations of this Earth system picture). State-of-the-art evidence garnered from mass extinctions in the Phanerozoic indicates that a couple of them were characterized by temperature changes of $\sim 10$ K (e.g., Dal Corso et al., 2022; Morgan et al., 2022). Therefore, upon adopting a fiducial temperature of $T_{0}=T_{\oplus}\equiv 288$ K for the planet and heating of $\Delta T\approx 6$ K, we duly obtain $\Delta T/T_{0}\approx 0.02$.
Second, we pivot to a relatively general phenomenon grounded in physics, since it may be argued that the temperature increases delineated in the preceding paragraphs are innately parochial, to wit, overly reliant on Earth’s biomes and biota. A number of numerical studies have demonstrated that the “moist greenhouse effect” can be instantiated at a typical temperature of around $330\pm 10$ K (Kasting et al., 1993; Leconte et al., 2013; Gómez-Leal et al., 2018; Ramirez, 2020), although the precise value is predicted to be modulated by a variety of stellar and planetary parameters (e.g., Wolf and Toon, 2015; Popp et al., 2016; Fujii et al., 2017; Kopparapu et al., 2017).
As a moist greenhouse effect could result in the depletion of substantial inventories of liquid water, it represents a plausible candidate for stymieing planetary habitability for life-as-we-know-it. For the fiducial temperature of $T_{0}=288$ K chosen earlier, a moist greenhouse threshold of approximately $330$ K amounts to setting $\Delta T\approx 42$ K. In turn, the fractional temperature increase in this scenario translates to $\Delta T/T_{0}\approx 0.15$, which was introduced previously.
One general point that should be recognized is that $\Delta T/T_{0}$ is exclusively a function of the ratio of the heating $Q_{\rm tech}$ and the stellar flux at the surface $S\left(1-A\right)/4$, as seen from (7). Thus, it does not incorporate additional heating that might arise from climate changes (e.g., atmospheric composition) during the growth in energy consumption. Hence, it seems prudent to regard $\Delta T/T_{0}$ calculated from (7) as an underestimate of the actual increase in the surface temperature, because of which the attendant timescale gauged herein to reach a given $\Delta T/T_{0}$ may represent an overestimate (i.e., upper bound of sorts).
2.6 Maximum lifespan of technospheres
Figure 3: The upper bound on the lifespan of technospheres with increasing energy usage, computed as the maximum time the planet spends within the habitable zone of a star of effective temperature $T_{\rm eff}$; it is assumed that the power consumption of technospheres displays exponential growth at a constant annual rate of $a$. The continuous (dotted) lines are based on the conservative (optimistic) HZ boundaries computed in Kopparapu et al. (2013).
The circumstellar habitable zone (HZ) is a cornerstone of astrobiology, with roots dating back to the 19th century (Lingam, 2021). The lifetime of the continuous habitable zone – which solely involves abiotic (i.e., neither biological nor technological) factors – was analyzed in several publications (Kasting et al., 1993; Kopparapu et al., 2013; Rushby et al., 2013; Gallet et al., 2017; Levenson, 2021; Ware et al., 2022; Tuchow and Wright, 2023), based on the consideration of a gradually increasing stellar luminosity. We will now proceed to estimate the upper bound on the lifetime of technospheres, where the key determinant of longevity is the energy consumption growth rate of technological species.
In classic models of the HZ, the inner and outer limits are expressed in terms of the corresponding top-of-atmosphere stellar flux $S_{\rm max}$ and $S_{\rm min}$, respectively regulated by the maximum and moist greenhouse effect (Kasting et al., 1993; Kopparapu et al., 2013; Ramirez, 2018; Ramirez et al., 2019; Lingam and Balbi, 2024). These thresholds can, to a certain degree, be perceived as “universal” because they are built from well-defined atmospheric and surface physics and chemistry, and hence they may pose fairly stringent constraints on planetary habitability, although exceptions could exist.
At an initial time (e.g., as when $P=P_{0}$ in Section 2.4), let us suppose that the planet harboring a technological species – albeit one that has not yet entered the exponential growth phase – is at the outer edge of the habitable zone and receives a stellar flux of $S_{\rm min}$. The planet cannot exist any further away from the host star, because liquid water would no longer be viable on the planetary surface (for a given atmospheric composition), and the world may therefore not be capable of sustaining a technological species with an Earth-like biochemistry (viz., entailing liquid water as the solvent).
With the onset of exponential growth in energy consumption, the surface temperature would rise due to the increasing additional heating resulting from technological activity, as revealed from (7). It may be valid to suppose that the stellar flux and the planetary albedo both remain constant throughout the period of interest (but we do briefly comment on exceptions to this implicit norm in Section 3), because this interval is relatively short, as demonstrated hereafter. Consequently, any changes in the net energy input at the surface are attributable to variations in $Q_{\rm tech}$ in our model.
We will now treat $Q_{\rm tech}$ as though it were an extra energy flux applied at the top of the atmosphere; the benefit will become apparent in the next paragraph. Of this energy input, only a fraction would pass through the atmosphere and reach the surface; if we denote the latter by $\tilde{Q}_{\rm tech}$, it is likely that $\tilde{Q}_{\rm tech}<Q_{\rm tech}$. Another way to frame this statement is that the real warming induced by $Q_{\rm tech}$ would be more severe than the assumption made herein, since this heat flux is actually engendered within the system (i.e., as an additional heat layer at the planetary surface). Thus, the lifetime of technospheres derived below constitutes an upper bound because the heating at the surface in our modified framework (which is $\tilde{Q}_{\rm tech}$) is smaller than the true heat flux $Q_{\rm tech}$.
Given that we have “lifted” $Q_{\rm tech}$ to the top of the atmosphere to calculate the upper bound on the technosphere lifetime, we can write the total top-of-the-atmosphere flux at any moment in time as
$S(t)\approx Q_{\rm tech}(t)+S_{\rm min},$ (19)
with the initial condition $S(t=0)\approx S_{\rm min}$, as motivated earlier. [^13] Strictly speaking, the initial condition should be $S(t=0)=S_{\rm min}+Q_{\rm tech}(t=0)$, but the second term on the RHS (i.e., waste heat flux) may be neglected compared to the first term. As $Q_{\rm tech}(t)$ grows exponentially with time, the same trend is evinced by $S(t)$. Eventually, the magnitude of $S(t)$ will exceed $S_{\rm max}$, at which stage the loss of liquid water and habitability is instantiated. In other words, because we require $S(t)\leq S_{\rm max}$ for maintaining habitable conditions, this relation is expressible as
$Q_{\rm tech}\leq S_{\rm max}-S_{\rm min}\equiv\Delta S,$ (20)
upon employing (19). Instead of choosing an initial stellar flux of $S_{\rm min}$, if we selected a higher value, then the upper bound for $Q_{\rm tech}$ analogous to (20) would decrease, as would the timescale to reach the threshold of $S_{\rm max}$. It is plausible, therefore, that the timescale(s) for truly crossing the boundaries of habitability may be lower than the upper bound computed herein.
The values of $S_{\rm max}$ and $S_{\rm min}$ exhibit some variation across publications (Shields et al., 2016; Lingam and Loeb, 2019), such as dependence on atmospheric composition (Pierrehumbert and Gaidos, 2011; Zsom, 2015; Ramirez, 2018; Madhusudhan et al., 2021), planetary mass and rotation rate (Kopparapu et al., 2014; Yang et al., 2014; Kopparapu et al., 2016), and the properties of biota (Schwieterman et al., 2019; Lingam, 2020), to name a handful. We shall adopt the well-known prescription of Kopparapu et al. (2013) for the HZ boundaries around F-, G-, K- and M-type stars. We will compute the duration $\Delta t$ required for $Q_{\rm tech}$ to surpass $\Delta S$, which represents the maximal lifetime of technospheres, as justified in the prior paragraphs. On invoking (16) and presuming that (18) holds true, we substitute these expressions in (20), thereby yielding
$\Delta t\approx\frac{\ln\left(\Delta S/Q_{\rm tech,0}\right)}{\ln\left(1+a\right)}\ {\rm years}$ (21)
where $Q_{\rm tech,0}\sim P_{0}/\left(4\pi R_{\oplus}^{2}\right)$ is the heating from technological activity at the initial time (i.e., commencement of exponential growth). This lifetime is shown in Figure 3 for various values of $a$, after drawing on both the conservative and optimistic HZ boundaries from Kopparapu et al. (2013). In all instances, we observe that the maximal lifetime of technospheres is $<1000$ years, provided that $a\gtrsim 0.01$ throughout the epoch of interest.
Lastly, a perusal of (20) reveals that the upper bound on $Q_{\rm tech}$ is set by $\Delta S$, from which we can duly estimate the maximal power $P_{\rm max}$ associated with the loss of habitability on an Earth-like planet containing a growing technological species. If (18) is applicable, we end up with $P_{\rm max}\sim 4\pi R_{\oplus}^{2}\Delta S$. Upon plugging this value in (15), the upper limit of the Kardashev scale ($K_{\rm max}$) attainable by such an ostensible technological species is found to be
$K_{\rm max}\approx\frac{1}{10}\log_{10}\left(\frac{4\pi R_{\oplus}^{2}\Delta S}{10^{6}\,\mathrm{W}}\right),$ (22)
and the ensuing results are plotted in Figure 4. It is evident that the maximal Kardashev scale $K_{\rm max}$ increases with the stellar temperature, but only by a minimal amount - in all the analyzed cases, $K_{\rm max}$ exists in a narrow range of about 0.850.850.850.850.90.90.90.9, slightly lower than unity. A corollary is that, for the simple model(s) in this paper, the maximal Kardashev scale under this particular set of assumptions may be less than unity. A vital facet of (22) is that this limit is, at first glimpse, independent of both the modality and rate (or equivalently timescale) of growth in energy consumption. Hence, this ceiling theoretically represents the maximal Kardashev level attainable by a technological species within a particular terrestrial planetary environment, irrespective of the specific energy growth model adopted, and might be accordingly construed as a physical (i.e., thermodynamical) limit endowed with a certain degree of generality.
Figure 4: The upper bound on the Kardashev scale level attainable by a technological species inhabiting an Earth-like planet in the habitable zone, before the loss of habitability due to the growth in energy consumption. The two scenarios depicted represent optimistic and conservative limits for staying within the habitable zone.
3 Discussion and Conclusions
As articulated in Section 1, waste heat production is an inevitable consequence of thermodynamics. In this paper, we investigated the myriad effects of technological species – and their accompanying technospheres – experiencing exponential growth with respect to their energy consumption and waste production.
Our analysis suggests that, if the energy growth rate is of order $1\%$ per year, the maximal lifetime of such putative technospheres is ephemeral compared to stellar evolution. From Figure 2, we notice that significant heating – which can drive the planet beyond potentially dangerous thresholds (consult Section 2.5) – due to exponential waste heat generation is predicted to occur on a short timescale of $\lesssim 1000$ years; these findings are broadly compatible with prior studies (von Hoerner, 1975; Mullan and Haqq-Misra, 2019). Likewise, a scrutiny of Figure 3 reveals that the upper bound on the lifetime of technospheres is relatively insensitive to stellar spectral type, and is merely hundreds of years in duration. These results stand in stark contrast to Earth’s long-lived biosphere that has existed for perhaps $\sim 4$ Gyr, as per the available geological evidence (Javaux, 2019; Lepot, 2020; Westall et al., 2023), phylogenetic inferences (Betts et al., 2018; Mahendrarajah et al., 2023; Moody et al., 2024), and simple mathematical models (Lingam and Loeb, 2017b).
We may identify three classes of trajectories that seem consistent with our modeling in this work, as listed below.
  1. 1.
    Technological species that pursue relentless exponential growth of energy consumption beyond the planet’s safe operating thresholds render themselves extinct on short timescales of typically $\lesssim 1000$ years.
  2. 2.
    Technological species transition from the phase of exponential growth in energy consumption to either an indefinite period of (near-)zero growth or even intervals of negative growth.
  3. 3.
    Technological species venture beyond their home planet(s), thereby utilizing space infrastructure for producing and dissipating energy, as well as for performing other technological activities.
We shall now tackle each of these trajectories in more detail, and briefly comment on their implications. We emphasize, however, that we do not claim to exhaust all conceivable routes in our précis. A comparatively exhaustive treatment of Earth’s technosphere $1000$ years from the present day, drawing on the domain of future studies, was recently explored in Haqq-Misra et al. (2024).
For instance, technological species may deliberately diminish the net incident stellar flux (which we held constant on the timescales considered) to offset increasing waste heat generation. [^14] On the other hand, such a reduction could translate to less energy accessible for, say, photovoltaic conversion, thus possibly stymieing growth (and dovetailing with Trajectory #2). Loosely overlapping with the forthcoming trajectory #3, this procedure may entail undertaking large-scale (macro)engineering activities near the surface and in the atmosphere – such as inputting aerosols to raise albedo (Rasch et al., 2008; Keith, 2013; MacMartin et al., 2018; Muri et al., 2018; MacMartin and Kravitz, 2019) – or even in outer space: constructing starshades (Early, 1989; Angel, 2006; Gaidos, 2017; Baum et al., 2022; Roy, 2022) and variants thereof (Pearson et al., 2006; Bewick et al., 2013; Borgue and Hein, 2023; Bromley et al., 2023) or nudging the planet into a wider orbit (Korycansky et al., 2001; McInnes, 2002; Korycansky, 2004)are a couple of such examples. While a subset of these projects have witnessed extensive theoretical explication and assessments, note that potential lacunae and issues with regard to the engineering, environmental, economic, social, political, and ethical dimensions have been broached (e.g., Vaughan and Lenton, 2011; Bellamy et al., 2012; Goldblatt and Watson, 2012; Caldeira et al., 2013; Irvine et al., 2016; Morton, 2016; Pamplany et al., 2020; Ricke et al., 2023).
With the aforementioned caveat and one credible example of a trajectory not elaborated in this paper out of the way, we will now proceed to elucidate the three categories previously introduced a couple of paragraphs ago.
Trajectory #1: If the majority of technological species are driven to extinction during their pursuit of exponential growth on account of violating one or more planetary thresholds, their mean lifetime during the detectable phase may be markedly low, as indicated earlier. In this specific scenario, the estimated number of communicative technological species in the Galaxy would drop commensurately as per the famous Drake equation formulated by Frank Drake (Drake, 1965), since one of the constituent factors is the average longevity of such species (Mamikunian and Briggs, 1965; Shklovskii and Sagan, 1966; Drake and Sobel, 1992; Burchell, 2006; Shostak, 2011; Vakoch and Dowd, 2015).
If the number of these technological species is indeed modest – based on the Drake equation and its probabilistic descendants (Glade et al., 2012; Maccone, 2012; Mieli et al., 2023), which are regulated by the longevity – this outcome might offer a (partial) “solution” to the famous Fermi paradox (Shermer, 2002; Webb, 2015; Cirkovic, 2018; Forgan, 2019): we have not encountered technological species because they are rare at any given moment in time. This explanation of the Fermi paradox in terms of long-term sustainability – or rather, the lack thereof – was surveyed by the likes of Dutil and Dumas (2009) and Haqq-Misra and Baum (2009), and fleshed out in subsequent publications by Frank et al. (2018); Mullan and Haqq-Misra (2019); Bologna and Aquino (2020); Lingam and Loeb (2021); Savitch et al. (2021); Wong and Bartlett (2022).
However, even if the majority of technological species are short-lived, in accordance with Lindy’s Law (Taleb, 2012), this facet does not necessarily preclude us from detecting their attendant technosignatures (Balbi and Grimaldi, 2024). As this trajectory encompasses energy-intensive species, a variety of detectable technosignatures are conceivable (e.g., Lingam and Loeb, 2021, Chapter 9), such as silicon-based photovoltaic cells and space-based mirrors for implementing stellar energy conversion (Korpela et al., 2015; Lingam and Loeb, 2017a; Kopparapu et al., 2024), nitrogen dioxide from combustion (Kopparapu et al., 2021), and urban heat islands (Kuhn and Berdyugina, 2015; Berdyugina et al., 2018). Moreover, a fraction of these signatures may persist over substantial timescales, even perhaps long after their creators have become extinct or departed the world (Holmes, 1991; Ćirković et al., 2019; Wright et al., 2022).
Trajectory #2: The pivotal outcomes that could stem from reifying protracted exponential growth of any given resource(s) have attracted intense debates since at least the classic treatise by Malthus (1798). A diverse collection of perspectives advanced by intellectuals spanning multifarious schools of thought – dating back to Mill (1848) at the minimum – have advocated in favor of humankind ultimately shifting to a paradigm of (near-)zero growth of that particular resource(s) sensu lato (i.e., a quasi-steady state in essence) or even embracing appropriate intervals of negative growth, which may be interpreted as degrowth (Georgescu-Roegen, 1971; Daly, 1973; Meadows et al., 1972; Hirsch, 1978; Daly, 1996; Burkett, 2006; Odum, 2007; Latouche, 2009; Kerschner, 2010; Foster et al., 2010; Bardi, 2011; Daly and Farley, 2011; Randers, 2012; D’Alisa et al., 2014; Costanza et al., 2015; Schmelzer, 2015; Schramski et al., 2015; Kallis et al., 2018; Jackson and Victor, 2019; Smil, 2019; Hickel, 2020; Kallis et al., 2020; Millward-Hopkins et al., 2020; Wiedmann et al., 2020; Herrington, 2021; Keyßer and Lenzen, 2021; McPhearson et al., 2021; Foster, 2022; Hickel et al., 2022; Saito, 2022; Schmelzer et al., 2022; Buch-Hansen et al., 2024). It should be recognized, however, that these paradigms have, in turn, garnered a plethora of criticisms, critiques, and responses (e.g., highlighting purported lacunae) from variegated standpoints (Ekins, 2000; Hart, 2005; Martínez-Alier et al., 2010; van den Bergh, 2011, 2017; Anderson and M’Gonigle, 2012; Fay, 2012; Hallegatte et al., 2012; Schwartzman, 2012; Naam, 2013; Quilley, 2013; Cosme et al., 2017; Kish and Quilley, 2017; Obama, 2017; Buch-Hansen, 2018; Liodakis, 2018; Strunz and Bartkowski, 2018; Büchs and Koch, 2019; Spash, 2020; Trainer, 2020; Durand et al., 2024; Gates, 2021; Polewsky et al., 2024; Ritchie, 2024).
For an annual growth rate ($a$) obeying $a\ll 1$, it can be shown that the maximal lifetime of technospheres is proportional to $1/a$ on taking this limit in (21). Hence, if a technological species is capable of drastically reducing the growth rate, its longevity will be proportionally enhanced, although the threshold imposed by (20) is still at play. By the same token, shifting from exponential growth to, say, power-law growth can also extend the lifespan of the technosphere because an exponential function increases faster than a power law in due course; this statement is likewise valid for logistic growth, which is tantamount to having the growth rate approach zero asymptotically. Lastly, the pursuit of growth might engender collapse, but not the extinction, of technological species, thence serving as a negative feedback suppressing growth and elevating their lifespan, albeit with a reduced energy footprint.
At this juncture, it is worth underscoring that, as long as some version of energy growth is existent, the magnitude of $Q_{\rm tech}$ would presumably continue to rise until the planetary thresholds from Section 2.5 are breached. In other words, these thresholds are typically applicable to generic planet-bound technological species, but the particular timescales for crossing them are contingent on the modalities of growth and the corrective actions taken by technological species. Since the latter are challenging to forecast, this work has focused on a single growth pattern (viz., exponential growth), with different ansatzen (e.g., non-monotonic patterns) representing the basis of desirable future work.
Thus, given that several studies have proposed that humanity is more likely to detect long-lived technological species (cf. Kipping et al., 2020; Balbi and Ćirković, 2021; Balbi and Grimaldi, 2024), it is not implausible that these species might have switched to a sustainable state of minimal (perhaps even near-zero or negative) energy growth or to slower modes (e.g., power laws) of growth; of the $10$ future scenarios envisioned for Earth’s technosphere in $1000$ years by Haqq-Misra et al. (2024), $5$ of them evince zero growth. One subtlety that should, however, be borne in mind is that stable technological species may entail modest energy consumption and waste heat dissipation, thereby perhaps lowering their likelihood of detection on the one hand while simultaneously boosting longevity on the other hand (cf. Tarter et al., 2018). [^15] https://www.kschroeder.com/weblog/the-deepening-paradox
Trajectory #3: Hitherto, we have chronicled either the looming specter of extinction because of the pursuit of untenable growth (Trajectory #1) or longevity with the possible stipulation of maintaining a restricted energy footprint (Trajectory #2), the latter of which stems from (20). It is natural to wonder, therefore, whether the existence of alternative pathways that can synthesize the pros of Trajectory #1 and Trajectory #2 is feasible from a theoretical viewpoint.
To resolve this matter, let us examine (20), which signifies an upper bound on the allowed heat flux. This limit was interpreted as capping the permitted energy consumption (i.e., power) and the Kardashev scale at the end of Section 2.6: refer to (22) and the associated exposition. However, this line of reasoning is implicitly valid only if the total surface area is held fixed. In contrast, if technological species were to expand their activities into space – which has been advocated by many to yield a myriad of benefits to humankind (Bernal, 1929; Stapledon, 1937; Dyson, 1960; O’Neill, 1977; Finney and Jones, 1985; Sagan, 1994; Bostrom, 2003; Harrison, 2001; Dick and Launius, 2007; Beech, 2009; Crawford, 2010; Rovetto, 2013; Munévar, 2014; Friedman, 2015; Crawford, 2016; Shelhamer, 2017; Zubrin, 2019, 2021; Ćirković, 2022; Johnson and Roy, 2023), although this stance has concomitantly garnered its share of critiques (Williams, 2010; Valentine, 2012; Slobodian, 2015; Schwartz and Milligan, 2016; Billings, 2019; Szocik, 2019; Traphagan, 2019; Deudney, 2020; Kovic, 2021; Gunderson et al., 2021; Schwartz et al., 2021) – these entities could decide to deploy space infrastructure for generating energy and dissipating heat over a much wider area (cf. Zotin et al., 2001; Wong, 2017). [^16] From the perspective of creating heat sinks, a distinctly speculative example by Opatrný et al. (2017) proposed that technological species with $K\gtrsim 1$ might control and utilize small black holes for this purpose (see also Loeb 2024). By doing so, the aforementioned bound on the power and the Kardashev scale may be circumvented to some degree.
Technological species belonging to Trajectory #3 might, in principle, readily surpass a Kardashev scale of unity – unlike their planet-based counterparts in Section 2.6 exhibiting $K_{\rm max}\approx 1$ – without perhaps sacrificing their longevity or sustainability (we address this theme in the next two paragraphs); the scenario S9 (out of $10$) outlined in Haqq-Misra et al. (2024) for Earth’s technosphere $1000$ years in the future envisages expansion into space by postbiological agents (artificial intelligences) unfolding at a growth rate of $1\%$. Trajectory #3, as remarked before in Section 3, shares select thematic links (albeit not equivalence) with alternative pursuits of large-scale engineering on the planet or in space as potential solutions. In case such technological species are viable and actually exist in the Milky Way, they may present us with a relatively strong opportunity to detect their accompanying technosignatures; the possible salient characteristics (e.g., optimization) of this category of technological species were discussed, inter alia, by Ćirković and Bradbury (2006).
We caution that unceasing exponential growth, even amidst the much broader horizons of space, might be rendered impractical at some stage in particular cases (as touched on shortly): if so, technological species may eventually adopt a variant of Trajectory #2 and curb their growth. Furthermore, whether Trajectory #3 can be instantiated while conceivably grappling with the specter of Trajectory #1 at a certain juncture (e.g., early on in history) could constitute an open question. An interesting feature of partial relevance in this context is the timescales corresponding to various phenomena. We have shown in this paper that the exponential growth of putative technological species on Earth-like planets can become unsustainable in $\sim 10^{3}$ yr under specific assumptions; whereas the duration for effectuating self-sustaining and stable planetary-scale modification (terraforming) of another world in toto might be up to $\lesssim 10^{5}$ yr with current human capabilities (cf. McKay and Marinova, 2001; Beech, 2009), for reaching neighboring stars with chemical propulsion is conventionally $\sim 10^{5}$ yr, and for settling the entire Milky Way is estimated to be $\sim 10^{7}$$10^{9}$ yr (Jones, 1976, 1981; Cotta and Morales, 2009; Nicholson and Forgan, 2013; Lingam, 2016; Carroll-Nellenback et al., 2019; Haqq-Misra and Fauchez, 2022).
We will not delve further into the complex dynamics of space settlement and Galactic expansion because this is not the central theme of the paper, and equally crucially, there are a multitude of intricate variables at play, ranging from engineering (e.g., propulsion systems, self-replication) and scientific to sociological and economic considerations. It has been suggested by some authors that exponential growth of technological species on a Galactic scale is ultimately not viable, due to bottlenecks on availability of resources or energy (von Hoerner, 1975; Newman and Sagan, 1981; Mullan and Haqq-Misra, 2019).
Coda: To round off this discourse, two of the three core themes underpinning astrobiology (arguably comprising its raison d’être) are centered on the fundamental questions of: (1) the putative existence and distribution of extraterrestrial life; and (2) the future of life on Earth and beyond (e.g., Chyba and Hand, 2005; Des Marais et al., 2008; Domagal-Goldman et al., 2016; Lingam and Balbi, 2024). By employing simple thermodynamic models, we have attempted to gauge the maximal lifetime of putative technological species and their technospheres characterized by exponential growth of energy consumption (a category which crucially encompasses humans in this epoch), since this endeavor has a direct bearing on the preceding two themes in this discipline.
Hence, even while several avenues that evidently warrant further research beyond this preliminary undertaking are apparent – such as notably harnessing sophisticated climate models to self-consistently calculate the temporal climate (co)evolution of temperate terrestrial planets hosting technospheres fuelled by sundry sources of energy or incorporating alternative modes of energy growth (e.g., logistic or power law growth) – which might consequently refine or modify (or perhaps even contradict) our central findings, we nevertheless hope that this work may serve to capture some of the salient physical details, and pave the way for in-depth treatments to chart the trajectories accessible by technological species in the Universe, including humans, as well as the ensuing signatures of technological activities (viz., technosignatures).
Acknowledgments
We are grateful to the reviewers for their constructive reports, which helped improve the quality and scope of the paper.